Please graph both graphs.

The fill in the blanks for the written are in order:
(0, 0.47, 2.14, 45.01)
(Elastic, inelastic, unit elastic)
(A decrease, an increase, no change)
(Elastic, inelastic, unit elastic)
Use the green rectangle (triangle symbols) to compute total revenue at various prices along the demand curve. Note: You will not be graded on any changes made to this graph. 300 275 250 225 200 175 150 125 100 75 50 Total Revenue 0 612 18 24 30 36 42 48 54 60 66 72 QUANTITY (Bikes)
media%2F8dd%2F8dda1461-799c-421b-a732-ed
media%2Fbe4%2Fbe41e00c-9728-4d23-b36a-e0
Answer #1
P ($) Q TR ($)
50 54 2700
75 48 3600
100 42 4200
125 36 4500
150 30 4500
175 24 4200
200 18 3600

Total revenue = P x Q

Total Revenue (S) 5000 4500 4000 3500 g 3000 2500 2000 1500 1000 500 Total Revenue (S) 50 100 150 200 250 Price ($)

PE = (P1 + P2)/(Q1 + Q2) x (Q2 - Q1)/(P2 - P1)

The price elasticity of demand between points A and B is approximately = (24 - 18)/(175 - 200) x 375/42 = 2.148

Because the demand between points A and B is elastic, a $25 per bike increase in price will lead to decrease in total revenue per day.

In general, in order for a price decrease to cause an increase in total revenue, demand must be elastic. (quantity demanded should change by a greater percentage in comparison to change in price percentage)